kayla is twice as old as her sister, jenna. four years ago kayla was 3 times jennas age. how old is jenna now?
Answer by Cromlix(4381)

kayla is twice as old as her sister, jenna.
K = 2J
four years ago kayla was 3 times jenna's age
K - 4 = 3(J - 4)
Replace K with 2J.
2J - 4 = 3(J - 4)
2J - 4 = 3J - 12
Collect like terms.
2J - 3J = - 12 + 4
-J = -8
J = 8
Jenna is 8 years old
Kayla is 16 years old.
Hope this helps.
:-)